OpenStudy (anonymous):

If f(z)=(z+1)/(z-1), then find f^1991(2+i)

OpenStudy (anonymous):

Full solution will get a medal (please tell me how-first time user)

OpenStudy (anonymous):

I got (2-i)^1991 and then what?

OpenStudy (mathmath333):

\(\large\tt \color{black}{ f^1991(2+i)=(\dfrac{2+i+1}{2+i-1})^{1991}}\) \(\large\tt \color{black}{ =(\dfrac{3+i}{1+i})^{1991}}\) \(\large\tt \color{black}{ =(\dfrac{3+i}{1+i}\times \dfrac{1-i}{1-i})^{1991}}\) \(\large\tt \color{black}{=(2-i)^{1991}}\) \(\large\tt \color{black}{=(2-i)\times(2-i)^{1990}}\) \(\large\tt \color{black}{=(2-i)\times(2-i)^{2(945)}}\) \(\large\tt \color{black}{=(2-i)\times(2^2-2i+i^2)^{945}}\) \(\large\tt \color{black}{=(2-i)\times(3-i)^{2(945)}}\) \(\large\tt \color{black}{=(2-i)(3-i)(3-1)^{2(472)}}\) like this u have to solve it untill u get the answer

OpenStudy (mathmath333):

second last term is incorrect ,it should be this \(\large\tt \color{black}{=(2-i)\times(3-i)^{945}}\)

OpenStudy (anonymous):

but then it would go on infinitely

OpenStudy (mathmath333):

yes i think so